# A Unified Approach for Multi-Granularity Search over Spatial Datasets

### v3

@@ Abstract.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For™ any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
@@ 1. Introduction
Consider a scenario where traffic analysts want to analyze traffic patterns in a metropolitan area. Firstly, they can define a query range of $R$ (auctus, ; Zacharatou2019, ; BrinkhoffKS93, ) to locate all relevant traffic datasets within that area for analysis, as illustrated on the left side of Figure1(a). Then, to identify the most efficient commuting routes, the analysts can input a specific traffic dataset $Q$ and search for datasets similar to $Q$ based on defined similarity measures (YangWZ2022, ; Wang2017, ; Wangsheng2018, ) for data analysis, illustrated on the right side of Figure 1(a). This search manner is known as exemplar search (Rezig2021, ; YewLML23, ; WangMGBCDSW24, ; Li0SWYY24, ; Mottin2014a, ; YangZ0H0020, ). Furthermore, if the analysts want to find points of interest, such as gas stations or rest areas along the searched routes $D$, they can specify a query range $R$ along with the relevant traffic dataset to retrieve these points, as depicted on the left side of Fig.1(b)(foursquare, ; Wang2019d, ; ZHANGSongnian2021, ; Qi2020, ). Additionally, on the right side of Figure1(b), the analysts can identify the nearest gas stations for each location in the dataset $Q$ (Qi2020, ; ZHANGSongnian2021, ; Dong2019b, ; ChoudhuryCSC16, ; CorralMTV2000, ), which helps commuters find convenient stops along their routes (YinWWCZ17, ; yao2016poi, ).
@@ 2.2. Spatial Data Point Search
$k$ Nearest Neighbor search.  The $k$NN (YuCWS07, ; ChenCCT15, ; Qi2020, ; G2010, ; NutanongS13, ) is defined as finding $k$ points in a given set that are closest to a given query point. The $k$NN queries have been extensively studied in the literature. To find the $k$ closest neighbor for any given query point, we can utilize the fundamental $k$NN search approach, which is often called the brute force (BF) method. For the given query points, it scans the entire dataset to find the $k$ closest points based on the distances between the query point and all other data points, which is computationally intensive. Thus, to efficiently accelerate the $k$NN search, a variety of indexes with different heuristic improvements have been proposed, such as R tree (Guttman84, ), K-Dimensional tree (KD-tree) (Sproull91, ), and Ball-tree (Omohundro1989, ). While these indexes can be implemented to find $k$ nearest neighbor points for a given point, they cannot support both spatial dataset search and spatial data point search, which is the focus of this work.
@@
3. Definitions
In this section, we first introduce the data model and multiple similarity metrics between spatial datasets. Subsequently, we introduce coarse-grained dataset search and fine-grained data point search, followed by the formalization of the problem of multi-granularity spatial data search.
@@ 3.1. Data Model
(Spatial Data Point). A spatial data point $p=(x,y,v_{1},\cdots,$ $v_{d-2})$ is a d-dimension vector ($d\geq 2$), which has at least a latitude $x$ and a longitude $y$, and can have $d-2$ other attributes.
@@
[…]
@@
Each cell, denoted as $c$, can be treated as an element. The coordinates of cell $c$ can be converted into a unique non-negative integer $c.id$ by interleaving the binary representations of its coordinate values (Peng2016, ; Zacharatou2019, ; SieranojaF18, ; YangWZ2022, ). This transformation yields consecutive IDs in the range $[0,2^{\theta}\times 2^{\theta}-1]$. In this way, a spatial dataset $D$ can be represented as a sorted integer set consisting of a sequence of cell IDs. The formal definition of z-order signature is as follows.
@@ 3.2.1. Overlap-based Similarity Measures
(Intersecting Area (IA)) Given a query dataset $Q$ and a spatial dataset $D$, the IA between $Q$ and $D$ is the size of the intersecting area of two given datasets’ MBR, i.e., IA$(Q,D)=\prod_{i=1}^{2}l(i,Q,D),$ where $l(i,Q,D)$ denotes the intersecting length of two MBRs in $i$-th dimension (Zomorodian2000, ).
@@ 3.3. Types of Dataset Search
(Top-$k$ Exemplar Dataset Search (ExempS)). Given a query dataset $Q$ and a spatial data repository $\mathcal{D}$, a top-$k$ exemplar search returns $k$ datasets from $\mathcal{D}$ that are most similar to $Q$, i.e., it retrieves $\mathcal{D}_{q}\subseteq\mathcal{D}$ with $k$ datasets such that: $\forall D\in\mathcal{D}_{q}$ and $\forall D^{\prime}\in\mathcal{D}\backslash\mathcal{D}_{q}$, we have $dist(Q,D)\leq dist(Q,D^{\prime})$, where $dist$ is any of the distance measures defined in Section 3.2.
@@ 3.4. Types of Data Point Search
(Range-based Point Search (RangeP)). Given a query range ${R}$, and a dataset $D\in\mathcal{D}_{q}$, the range-based point search returns all the points of $D$ inside the range $R$.
@@ 1 List of dataset root nodes $L\leftarrow\emptyset$;
       $N_{D_{i}}\leftarrow$SplitSpace($D_{i}$,$f$,$d$,$\phi$); $\triangleright$ bottom-level index
@@
$N_{root}\leftarrow\textsf{SplitSpace}(L,f,d,\emptyset$)$\triangleright$ upper-level index
@@ 23                   $d_{split}\leftarrow i$;
36      $N_{s}.N_{l}\leftarrow$ SplitSpace($leftChildList$, $f$, $d$, $\phi$);
37       $N_{s}.N_{r}\leftarrow$ SplitSpace($rightChildList$, $f$, $d$, $\phi$);
@@ 46      $r{{}^{\prime}}\leftarrow\phi[pos-1]$;
48             RefineBottomUp$(N_{D_{i}},r^{\prime})$;
49   Function RefineBottomUp($N$, $r{{}^{\prime}}$):
       Input: $N$: dataset index node, $r{{}^{\prime}}$: radius threshold
50       if $N$ is a leaf node then
@@ 5.1. Bottom Level: Index of Each Single Dataset
Geometrically, the gap $g_{i}$ for a radius $\phi[i]$ computes the distance between the point ($i$, $\phi[i]$) and its projection on the segment connecting the first point (00, $\phi[0]$) and the last point ($|\phi|-1$, $\phi[|\phi|-1]$), as shown in Figure 6(b). The radius with the largest gap is selected as the threshold $r^{\prime}$ (see Lines 1 to 1), as it is the turning point where most leaf nodes start to have a similar radius. Thus, leaf nodes with a radius greater than $r^{\prime}$ are refined by removing potential outliers.
@@ 6.1.1. Accelerating Range-based Dataset Search
Input: $Q$ or $R$: query dataset or range, $\mathcal{D}$: data repository, $N$: a node in the index, $k$: number of results
@@ 1 $\mathcal{R}esult\leftarrow\emptyset$;
2 if $N$ is a dataset root node then
@@
5            Add ${N.id}$ and update $\mathcal{R}esult$;
@@ 8                   $dis$leftarrow HausCompute($Q$, $N$);
11                         $dis$leftarrow Compute $dis$ between $Q$ and $N$;
@@
14                  Add ${N.id}$ and rank $\mathcal{R}esult$ increasingly by $dis$;
@@
21                   $dis\leftarrow$ compute $dis$ between $Q$ and $N_{c}$;
@@ 6.1.2. Accelerating Top-$k$ Exemplar Dataset Search
Output: $UB$: Hausdorff distance between $Q$ and $D$
@@ 1 $N_{Q}\leftarrow$ create a dataset root node of $Q$;
3       $^{\uparrow}\leftarrow\textsf{CreateAscendingPQ}()$PQ ( ), $PQ^{\uparrow}.Insert(N_{D},0)$;
[…]
@@
7      $[N_{Q},UB,PQ^{\uparrow}]\leftarrow PQ^{\downarrow}.Dequeue()$;
8      $[N_{D},LB]\leftarrow PQ^{\uparrow}.Head()$, $terminate\leftarrow false$;
@@ 29   Function TraversalDIndex($N_{D},PQ^{\uparrow},PQ^{\downarrow},UB$):
       Input: $N_{D}$: dataset node, $PQ^{\uparrow}$: main queue, $PQ^{\downarrow}$: second queue, $UB$: the upper bound
@@
31      $[N_{D},LB]\leftarrow PQ^{\uparrow}.Dequene()$, $minUB\leftarrow UB$;
@@ 7.1. Settings
  • Chicago taxi trips (chicago, ) dataset, which has 194,992,061 trips in total. As all the trips are quantized every 15 minutes, we split the whole dataset every 15 minutes to create a set of sub-datasets, and there are 283,375 datasets in total;
@@
  • $R$: It determines the range of query rectangular, where $\epsilon$ is the value obtained when the resolution $\theta$ is the default value $5$ according to the Equation 8 in Section 6.1.2.
@@
Methods for Comparison.  We compare our proposed Spadas system against the following baselines.
  • RangeS (Def. 8): our range-based dataset search supported by the Spadas system.
  • IA (Def. 5): our top-$k$ exemplar dataset search based on the intersecting area supported by the Spadas system.
  • GBO (Def. 6): our top-$k$ exemplar dataset search based on the grid overlap supported by the Spadas system.
  • ExactHaus: our top-$k$ exemplar dataset search based on the exact Hausdorff distance with a fast bound estimation technique.
  • ApproHaus: our top-$k$ exemplar dataset search based on the approximate Hausdorff distance with the error-bounded approximation technique.
  • RangeP (Def. 10): our range-based data point search supported by the Spadas system.
  • NNP (Def. 11): our nearest neighbor point search supported by the Spadas system.
@@
  • ScanHaus (Adelfio2011, ): It designs the lower and upper bounds of Hausdorff by using two point sets’ MBR, and combines branch-and-bound to accelerate filter out dissimilar datasets.
@@ 7.2. An Overview of spadas
Figure 10. Index’s construction time and occupied space with increasing data repository scale $m=\{10^{-3},10^{-2},10^{-1}$, 1}×$\times|\mathcal{D}|$.
@@ 7.3.2. Top-$k$ Point-wise Query
(b) Runtime of Top-$k$Hausdorff Search
@@
Effect of Dimension $d$.  As the Chicago dataset’s dimension is the highest, we use it to verify the performance of three representative operations in Figure 21. Firstly, from the left of Figure 21, we can see that the query time of the two overlap-based methods does not change much with the increase of dimension, and IA is much faster than GBO. The right of Figure 21 shows the search time of the top-$k$ Hausdorff. The experimental result shows that with the increase of $d$, the running time of the ScanHaus and IncHaus increases, while the times of the ExactHaus do not change significantly. This indicates that dimension significantly affects the effectiveness of calculating distances between MBRs.
@@ 7.4. Efficiency of Data Point Search
Nearest Neighbor Point Search.  We mainly compare our NNP search operation presented in Section 6.2 against the k-nearest neighbor (kNN(Taha2015, ) that works similarly to the Hausdorff distance (Nutanong2011, ). Figure 23 shows that our method can accelerate the NNP operation significantly, which achieves up to seven orders of magnitude speedup over kNN. This is because the kNN shares a common procedure with NNP to find the nearest points for points in the query datasets, and we can save much time by continuously traversing the unified index.
@@ 7.5. Case Study
Fig. 23 shows the visualization of the results that conduct trajectory search and analysis based on the Spadas system developed in this paper. Specifically, we added a transportation dataset collection to Spadas that includes a variety of trajectory datasets from Washington, D.C. An analyst aims to find $k$ trajectory datasets that are closely similar to a given query dataset from the Spadas to perform trajectory analysis tasks, such as trajectory near-duplicate detection (XiaoWLYW11, ) and traffic congestion prediction (ZhengZYSZ14, ; ZhouTWN13, ).
As shown in Fig. 23(a), the analyst first inputs a query dataset $Q$ from Washington, D.C. into Spadas, aiming to find the top 5 datasets that are similar to $Q$. The results found based on IA, GBO, and Haus similarity measures are shown in Figures 23(b), (c), and (d). We can see that the result datasets obtained by IA are coarse-grained compared with GBO and Haus. For example, although the intersecting area between the result dataset $D_{5}$ and $Q$ is quite large, their data point distributions are not similar. In contrast, the results obtained by Haus and GBO are more similar to the query $Q$, but Haus offers a more fine-grained similarity. For example, the result $D_{2}$ obtained by Haus is more similar to $Q$ in terms of dataset distribution than $D_{2}$ obtained by GBO. The analyst can use these retrieved datasets, along with $Q$, for trajectory data analysis tasks.

### v5

@@ Abstract.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For™ any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
@@ 1. Introduction
Consider a scenario where traffic analysts want to analyze traffic patterns in a metropolitan area. Firstly, they can define a query range of $R$ (auctus, ; Zacharatou2019, ; BrinkhoffKS93, ) to locate all relevant traffic datasets within that area for analysis, as illustrated on the left side of Figure1(a). Then, to identify the most efficient commuting routes, the analysts can input a specific traffic dataset $Q$ and search for datasets similar to $Q$ based on defined similarity measures (YangWZ2022, ; Wang2017, ; Wangsheng2018, ) for data analysis, illustrated on the right side of Figure 1(a). This search manner is known as exemplar search (Rezig2021, ; YewLML23, ; WangMGBCDSW24, ; Li0SWYY24, ; Mottin2014a, ; YangZ0H0020, ). Furthermore, if the analysts want to find points of interest, such as gas stations or rest areas along the searched routes $D$, they can specify a query range $R$ along with the relevant traffic dataset to retrieve these points, as depicted on the left side of Figure1(b)(foursquare, ; Wang2019d, ; ZHANGSongnian2021, ; Qi2020, ). Additionally, on the right side of Figure1(b), the analysts can identify the nearest gas stations for each location in the dataset $Q$ (Qi2020, ; ZHANGSongnian2021, ; Dong2019b, ; ChoudhuryCSC16, ; CorralMTV2000, ), which helps commuters find convenient stops along their routes (YinWWCZ17, ; yao2016poi, ).
@@ 2.2. Spatial Data Point Search
$k$ Nearest Neighbor search.  The kkNN (YuCWS07, ; ChenCCT15, ; Qi2020, ; G2010, ; NutanongS13, ) is defined as finding $k$ points in a given set that are closest to a given query point. The kkNN queries have been extensively studied in the literature. To find the $k$ closest neighbor for any given query point, we can utilize the fundamental kkNN search approach, which is often called the brute force (BF) method. For the given query points, it scans the entire dataset to find the $k$ closest points based on the distances between the query point and all other data points, which is computationally intensive. Thus, to efficiently accelerate the kkNN search, a variety of indexes with different heuristic improvements have been proposed, such as R tree (Guttman84, ), K-Dimensional tree (KD-tree) (Sproull91, ), and Ball-tree (Omohundro1989, ). While these indexes can be implemented to find $k$ nearest neighbor points for a given point, they cannot support both spatial dataset search and spatial data point search, which is the focus of this work.
@@
3. Preliminaries
In this section, we first introduce the data model and multiple similarity metrics between spatial datasets. Subsequently, we introduce coarse-grained dataset search and fine-grained data point search, followed by the formalization of the problem of multi-granularity spatial data search. Frequently used notations are summarized in Appendix 9.1.
@@ 3.1. Data Model
(Spatial Data Point). A spatial data point $p=(x,y,v_{1},\cdots,$ $v_{d-2})$ is a d-dimension vector ($d\geq 2$), which has at least a latitude $x$ and a longitude $y$, and can have d2d-2 other attributes.
@@
(Spatial Dataset). A spatial dataset $D$ contains a set of $2$-dimension points, $i.e.,D=\{p_{1},p_{2},\cdots,p_{|D|}\}$, where $|D|$ is the size of $D$. The minimum bounding rectangle (MBR) of $D$, whose sides are parallel to the $x$ and $y$ axes and minimally enclose all points, is specified by the bottom left corner $b^{\downarrow}$ and upper right corner $b^{\uparrow}$, i.e.,
@@
Each cell, denoted as $c$, can be treated as an element. The coordinates of cell $c$ can be converted into a unique non-negative integer $c.id$ by interleaving the binary representations of its coordinate values (Peng2016, ; Zacharatou2019, ; SieranojaF18, ; YangWZ2022, ). This transformation yields consecutive IDs in the range $0,2^{\theta}\times 2^{\theta}-1]$. In this way, a spatial dataset $D$ can be represented as a sorted integer set consisting of a sequence of cell IDs. The formal definition of z-order signature is as follows.
@@ 3.2.1. Overlap-based Similarity Measures
[…]
@@ 3.3. Types of Dataset Search
(Top-$k$ Exemplar Dataset Search (ExempS)). Given a query dataset $Q$ and a spatial data repository $\mathcal{D}$, a top-$k$ exemplar search returns $k$ datasets from $\mathcal{D}$ that are most similar to $Q$, i.e., it retrieves $\mathcal{D}_{q}\subseteq\mathcal{D}$ with $k$ datasets such that: $\forall D\in\mathcal{D}_{q}$ and $\forall D^{\prime}\in\mathcal{D}\backslash\mathcal{D}_{q}$, we have $dist(Q,D)\leq dist(Q,D^{\prime})$, where distdist is any of the distance measures defined in Section 3.2.
@@ 3.4. Types of Data Point Search
(Range-based Point Search (RangeP)). Given a query range R{R}, and a dataset $D\in\mathcal{D}_{q}$, the range-based point search returns all the points of $D$ inside the range $R$.
@@ 1 List of dataset root nodes $L\leftarrow\emptyset$;
[…]
@@
[…]
@@ 23        $d_{split}\leftarrow i$;
36 $N_{s}.N_{l}\leftarrow$ SplitSpace(leftChildListleftChildList, $f$, $d$, $\phi$);
37  $N_{s}.N_{r}\leftarrow$ SplitSpace(rightChildListrightChildList, $f$, $d$, $\phi$);
@@ 47 $r{{}^{\prime}}\leftarrow\phi[pos-1]$;
[…]
50   Function RefineBottomUp(NN, $r{{}^{\prime}}$):
  Input: NN: dataset index node, $r{{}^{\prime}}$: radius threshold
51  if NN is a leaf node then
@@ 5.1. Bottom Level: Index of Each Single Dataset
Geometrically, the gap $g_{i}$ for a radius $\phi[i]$ computes the distance between the point ($i$, $\phi[i]$) and its projection on the segment connecting the first point (0, $\phi[0]$) and the last point ($|\phi|-1$, $\phi[|\phi|-1]$), as shown in Figure 6(b). The radius with the largest gap is selected as the threshold $r^{\prime}$ (see Lines 1 to 1), as it is the turning point where most leaf nodes start to have a similar radius. Thus, leaf nodes with a radius greater than $r^{\prime}$ are refined by removing potential outliers.
@@ 6.1.1. Accelerating Range-based Dataset Search
Input: $Q$ or $R$: query dataset or range, $\mathcal{D}$: data repository, NN: a node in the index, $k$: number of results
@@ 1 $\mathcal{R}esult\leftarrow\emptyset$;
2 if NN is a dataset root node then
@@
5    Add N.id{N.id} and update $\mathcal{R}esult$;
@@
11           disdis $\leftarrow$ Compute disdis between $Q$ and NN;
@@
14       Add N.id{N.id} and rank $\mathcal{R}esult$ increasingly by disdis;
@@
21        $dis\leftarrow$ compute disdis between $Q$ and $N_{c}$;
@@ 6.1.2. Accelerating Top-$k$ Exemplar Dataset Search
Output: UBUB: Hausdorff distance between $Q$ and $D$
@@ 1 $N_{Q}\leftarrow$ create a dataset root node of $Q$;
3  $PQ^{\uparrow}\leftarrow\textsf{CreateAscendingPQ}()$, $PQ^{\uparrow}.Insert(N_{D},0)$;
4 $PQ^{\downarrow}\leftarrow\textsf{CreateDescendingPQ}()$, $PQ^{\downarrow}.Insert(N_{Q},\infty,PQ^{\uparrow})$;
@@
7 $N_{Q},UB,PQ^{\uparrow}]\leftarrow PQ^{\downarrow}.Dequeue()$;
8 $N_{D},LB]\leftarrow PQ^{\uparrow}.Head()$, $terminate\leftarrow false$;
@@ 29   Function TraversalDIndex($N_{D},PQ^{\uparrow},PQ^{\downarrow},UB$):
  Input: $N_{D}$: dataset node, $PQ^{\uparrow}$: main queue, $PQ^{\downarrow}$: second queue, UBUB: the upper bound
@@
31 $N_{D},LB]\leftarrow PQ^{\uparrow}.Dequene()$, $minUB\leftarrow UB$;
@@ 7.1. Settings
  • Chicago taxi trips (chicago, ) dataset, which has 194,992,061 trips in total. As all the trips are quantized every 15 minutes, we split the whole dataset every 15 minutes to create a set of sub-datasets;
@@
  • $R$: It determines the range of query rectangular, where $\epsilon$ is the value obtained when the resolution $\theta$ is the default value 55 according to the Equation 8 in Section 6.1.2.
@@
Methods for Comparison.  We compare our proposed Spadas system against the following baselines. The first four algorithms assess overlap-based dataset search, the next six evaluate pair-wise dataset search, and the last three test data point search.
  • IA (Def. 5): our top-$k$ exemplar dataset search based on the intersecting area in Spadas.
  • GBO (Def. 6): our top-$k$ exemplar dataset search based on the grid overlap in Spadas.
  • RangeS (Def. 8): our range-based dataset search in Spadas.
@@
  • ExactHaus: our top-$k$ exemplar dataset search using exact Hausdorff distance with fast bound estimation.
  • ApproHaus: our top-$k$ exemplar dataset search using approximate Hausdorff distance with the error-bounded approximation.
  • ScanHaus (Adelfio2011, ): It designs the lower and upper bounds of Hausdorff by using two point sets’ MBR, and combines branch-and-bound to accelerate filtering out dissimilar datasets.
@@
  • RangeP (Def. 10): our range-based data point search in Spadas.
  • NNP (Def. 11): our nearest neighbor point search in Spadas.
@@ 7.2. An Overview of spadas
[…]
@@ 7.3.2. Top-$k$ Point-wise Query
(b) Runtime of Top-kkHausdorff Search
@@
Effect of Dimension $d$.  As the Chicago dataset’s dimension is the highest, we use it to verify the performance of three representative operations in Figure 21. Firstly, from the left of Figure 21, we can see that the query time of the two overlap-based methods does not change much with the increase of dimension, and IA is much faster than GBO. The right of Figure 21 shows the search time of the top-$k$ Hausdorff. The experimental result shows that with the increase of $d$, the running time of the ScanHaus and IncHaus increases, while the times of the ExactHaus do not change significantly. This indicates that dimension significantly affects the effectiveness of calculating distances between MBRs. In addition, to investigate the scalability of our proposed framework, we also conduct experiments in Appendix 9.3.
@@ 7.4. Efficiency of Data Point Search
Nearest Neighbor Point Search.  We mainly compare our NNP search operation presented in Section 6.2 against the k-nearest neighbor (kNN(Taha2015, ) that works similarly to the Hausdorff distance (Nutanong2011, ). Figure 23 shows that our method can accelerate the NNP operation significantly, which achieves up to seven orders of magnitude speedup over kNN. This is because the kNN shares a common procedure with NNP to find the nearest points for points in the query datasets. By continuously traversing the unified index, our method reduces computation time substantially.
@@ 7.5. Case Study
Figure 23 shows the visualization of the results that conduct trajectory search and analysis based on the Spadas system developed in this paper. Specifically, we added a transportation dataset collection to Spadas that includes a variety of trajectory datasets from Washington, D.C. An analyst aims to find $k$ trajectory datasets that are closely similar to a given query dataset from the Spadas to perform trajectory analysis tasks, such as trajectory near-duplicate detection (XiaoWLYW11, ) and traffic congestion prediction (ZhengZYSZ14, ; ZhouTWN13, ).
As shown in Figure 23(a), the analyst first inputs a query dataset $Q$ from Washington, D.C. into Spadas, aiming to find the top 5 datasets that are similar to $Q$. The results found based on IA, GBO, and Haus similarity measures are shown in Figures 23(b), (c), and (d). We can see that the result datasets obtained by IA are coarse-grained compared with GBO and Haus. For example, although the intersecting area between the result dataset $D_{5}$ and $Q$ is quite large, their data point distributions are not similar. In contrast, the results obtained by Haus and GBO are more similar to the query $Q$, but Haus offers a more fine-grained similarity. For example, the result $D_{2}$ obtained by Haus is more similar to $Q$ in terms of dataset distribution than $D_{2}$ obtained by GBO. The analyst can use these retrieved datasets, along with $Q$, for trajectory data analysis tasks.